40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Jul 20 11:58:45 2022, Paco, Summary, General, Jenne laser kaput? 
    Reply  Wed Jul 20 14:12:07 2022, Paco, Summary, General, Jenne laser kaput! 
       Reply  Wed Jul 20 15:58:52 2022, Koji, Summary, General, Jenne laser kaput! 
          Reply  Mon Jul 25 09:05:50 2022, Paco, Summary, General, Testing 950nm laser found in trash pile PXL_20220722_234600124.jpgPXL_20220722_234551918.jpg
             Reply  Fri Aug 5 17:03:31 2022, Yehonathan, Summary, General, Testing 950nm laser found in trash pile 20220805_164434.jpgLaserActuation_TF_Measurement.drawio.pdf
                Reply  Wed Sep 28 19:15:56 2022, Koji, Update, General, Testing 950nm laser found in trash pile PXL_20220929_013850989.MP.jpgPXL_20220929_013859439.jpgPXL_20220929_013911125.jpg
                   Reply  Wed Nov 30 18:27:07 2022, Paco, Update, General, Making the Jenne laser great again PXL_20221201_004623398.jpgPXL_20221201_015752792.jpg
Message ID: 17162     Entry time: Wed Sep 28 19:15:56 2022     In reply to: 17064     Reply to this: 17326
Author: Koji 
Type: Update 
Category: General 
Subject: Testing 950nm laser found in trash pile 

I don't know what was wrong with the past setup but the 950nm laser (QPHOTONICS QFLD-950-3S) just worked fine up to ~300MHz with basically the same setup.

A 20dB coupler picks up a small amount of the driving signal from the source signal of the network analyzer. This was fed to CHR. The fiber-coupled NewFocus PD RF output was connected to CHA.
The calibration of the response was done with the thru response (connect the source signal to the CHA via all the long cables).

Attachment 1 shows the response CHA/CHR. The output is somewhat flat up to 20MHz and goes down towards 100MHz, but still active up to 500MHz as long as the normalization with the New Focus PD works.
The structure around 200MHz~300MHz changes with how the wires of the clips are arranged. I have twisted and coiled them as shown and the notch disappeared. For the permanent setup we should keep the lines as short as possible and take care of the stray capacitance and the inductance.

Attachment 2 shows the setup at the network analyzer side. Nothing special.

Attachment 3 shows the setup at the laser side. The DB9 connector on the Jenne's laser has the negative output of the LD driver connected to the coax core and the positive output connected to the shield of the coax. Therefore the coax core (red clip) has to be connected to Pin 9 and the coax shield (black clip) to PIn 5.

Attachment 1: PXL_20220929_013850989.MP.jpg  4.711 MB  Uploaded Wed Sep 28 20:17:31 2022  | Hide | Hide all
PXL_20220929_013850989.MP.jpg
Attachment 2: PXL_20220929_013859439.jpg  2.581 MB  Uploaded Wed Sep 28 20:17:35 2022  | Hide | Hide all
PXL_20220929_013859439.jpg
Attachment 3: PXL_20220929_013911125.jpg  3.161 MB  Uploaded Wed Sep 28 20:17:39 2022  | Hide | Hide all
PXL_20220929_013911125.jpg
ELOG V3.1.3-